# thanks: $0$-0002-5209-872Xthanks: $0$-0002-1235-0505

### v1

@@ Abstract
The next generation of neutrino oscillation experiments, JUNO, DUNE, and HK, are under construction now and will collect data over the next decade and beyond. As there are no approved plans to follow up this program with more advanced neutrino oscillation experiments, we consider here one option that had gained considerable interest more than a decade ago: a neutrino factory. Such an experiment uses stored muons in a racetrack configuration with extremely well characterized decays reducing systematic uncertainties and providing for more oscillation channels. Such a machine could also be one step towards a high energy muon collider program. We consider a long-baseline configuration to SURF using the DUNE far detectors or modifications thereof, and compare the expected sensitivities to the three-flavor oscillation parameters to the anticipated results from DUNE and HK. We find that a neutrino factory can improve our understanding of CP violation and also aid in disentangling the complicated flavor puzzle.
I Introduction
@@
The next generation of neutrino oscillation experiments, including JUNO Abusleme et al. (2022), DUNE Abi et al. (2020), and HK Abe et al. (2015) are expected to make significant headway in addressing three of the remaining known unknowns in particle physics. Notably, they will have excellent $\gg 5\sigma$ sensitivity to the mass ordering (the sign of $\Delta m^{2}_{31}$), they will have good $\theta_{23}$ octant determination capabilities provided that it isn’t too close to maximal, and both DUNE and HK will each individually provide $>5\sigma$ sensitivity to CP violation for significant regions of parameter space. In addition, they will provide some resolution on measuring $\delta$ (we use the standard parameterization of the PMNS Pontecorvo (1957); Maki et al. (1962) matrix Denton and Pestes (2021)). Atmospheric neutrino experiments such as HK, the IceCube-upgrade Ishihara (2021), and KM3NeT-ORCA Adrian-Martinez et al. (2016) will also provide complementary information. For a recent overview of this picture, see Denton et al. (2022).
@@
While the outlook is hopefully rosy, it is important to carefully examine what, if anything, should come after these experiments in the neutrino oscillation program. Several experiments have been proposed with varying levels of detail. One is using a water-based liquid scintillator detector THEIA which combines the advantages of water and liquid scintillator, likely in place of one of the four LArTPC modules for DUNE Askins et al. (2020, 2022). Another is an additional HK-like tank in Korea experiencing a different portion of the same beam that HK experiences to target the second oscillation maximum in appearance Abe et al. (2018). There is also discussion about building a very large (much larger than HK) water Cherenkov detector in Sweden combined with additional beam upgrades of the ESS to create the ESSnuSB also to target the second oscillation maximum in Alekou et al. (2021, 2022); Abele et al. (2023).
@@
In this article we discuss a different option, a neutrino factory. Neutrino factories were seriously considered as an option to probe CP violation in the first decade of this millennium largely due to the expectation of small $\theta_{13}<1^{\circ}$ which severely limits the possible size of the CP signal De Rujula et al. (1999). In general, a neutrino factory is a muon storage ring in a racetrack configuration with long straight sections pointing towards the far detector and provides powerful beams of both muon and electron neutrinos of opposite charges Geer (1998). The initial spectrum of this decay is extremely well characterized unlike the initial spectrum from traditional fixed target neutrino oscillation experiments which primarily produce neutrinos from light meson decays with a range of energies. See Bogacz et al. (2022) for a recent review of the history of neutrino factories as well as a discussion of the non-oscillation physics program at such a machine.
@@
A neutrino factory may also make sense as a first step on the path to a multi-TeV high energy muon collider to probe the electroweak sector, see e.g. Accettura et al. (2023). This first step could be chronologically as a part of a technology demonstrator, a part of the accelerator chain, or both. In addition, the recent 2023 P5 report suggested consideration of neutrino physics to be done with a muon accelerator p5_ . A neutrino factory may also be a necessary successor to the upcoming experiments if there is a disagreement in the determination of the oscillation parameters among the different experiments; a situation that the neutrino field is well accustomed to.
@@
In this paper we will discuss the oscillation physics at a neutrino factory considering the current knowledge of the oscillation parameters and address how a modern look at the oscillation physics case for a neutrino factory compares to previous studies. We then derive numerical sensitivities for the precision on $\delta$ for a potential setup of a neutrino factory for two different baselines and energies, before we conclude.
II Neutrino Factory Setup
@@
  • The achievable maximal neutrino energy is higher at a neutrino factory as the produced muon neutrinos have energies up to the muon energy. Indeed the muon neutrino energy spectrum rises with energy instead of falling like at a fixed target neutrino oscillation experiment, see fig. 1.
  • The composition and the expected energy of the neutrino beam is well known as radiative corrections to muon decay are small Greub et al. (1994); Laing and Soler (2008); Tomalak (2022). Furthermore, no $\nu_{\tau}$ get produced in the source, allowing for clean searches for $\nu_{\tau}$ appearance channels.
  • The neutrino beam at a neutrino factory is composed of equally many neutrinos as anti-neutrinos, albeit of different flavors. In particular, the flux consists of 50% $\nu_{e}$ which allows for $\nu_{e}\to\nu_{\tau}$ searches, unlike at fixed target oscillation experiments where the $\nu_{e}$ flux is designed to be small.
  • The neutrino energy is tunable and flexible, which allows for physics studies at different, well-known energies.
The first studies of a neutrino factory studies were focused on measuring a small $\theta_{13}$, $\delta$, and the mass ordering and identified problems with degeneracies in the oscillation analysis Barger et al. (2002); Freund et al. (2001); Huber et al. (2006). For small values of $\theta_{13}$, charge ID (CID) of muons or electrons was crucial to suppress the beam background for the appearance channels and to measure $\delta$. For this reason, many studies assumed a 100 kT magnetized iron detector Choubey et al. (2011).
With the current and envisioned knowledge of all oscillation parameters and the present experimental landscape, the task of a future neutrino factory is no longer to discover the oscillation parameters but more likely to make precise measurements and/or potentially resolve any discrepancies identified in previous measurements. Therefore, with a different physics goal and a more modern experimental landscape, the experimental requirements on a neutrino factory also necessarily evolve.
@@
We consider two possible neutrino factory scenarios: the neutrino source is Fermilab and, to make use of existing facilities, the far detector to be located at SURF, resulting in a baseline of 1284.9 km. Alternatively, we also consider the neutrino source at Brookhaven National Laboratory, to make use of the AGS/RHIC/EIC accelerator complex, which leads to baseline of 2542.3 km to SURF. We assume as matter density for both configurations 2.848 g/$\text{cm}^{3}$. We assume that the far detector is based on a LArTPC configuration with a total fiducial target mass of 40 kT consistent with the envisioned DUNE detector. Based on previous exploratory studies of neutrino factories we assume $10^{21}$ muon (or anti-muon) decays per year Choubey et al. (2011).
@@
To simulate a neutrino factory we make use of DUNE’s far detector design Abi et al. (2021) in the GLoBES Huber et al. (2005) framework. We also consider oscillated tau neutrinos events with a tau neutrino CC cross section from Yaeggy (2023). We include tau events as background only Donini et al. (2011) and don’t consider tau signal events for this three-flavor oscillation study due to the large uncertainties on the signal efficiency at LAr detectors; including this channel would improve the sensitivity to measure the oscillation parameters, although likely at a subleading level due to the lower statistics because the $\nu_{\tau}$ cross section is suppressed at these energies. As the initial neutrino flux is well known we assume a normalization uncertainty of 2% on the electron signal and background, 5% on the muon signal and background, 10% on the NC background, 20% on the tau background.
III Sensitivities
@@
For our simulations we consider priors on the oscillation parameters other than $\delta$ from Gonzalez-Garcia et al. (2021) and we fix the MO to NO, as the MO as well as all oscillation parameters will be determined via multiple independent methods with future experiments and our results are largely independent of the true MO. Therefore we do not study the sensitivity of a neutrino factory on these parameters but instead study the possible improvements a neutrino factory can bring on the precision on $\delta$. This provides an estimate of the ability of a neutrino factory to differentiate among differing measurements from upcoming experiments.
We studied the sensitivity to the oscillation parameters as a function of muon energy as well as the ratio of muon to anti-muon running time. For the FNAL-SURF (BNL-SURF) baseline we find that a muon energy of $E_{\mu}=5$ (8) GeV with equal muon and anti-muon running time allows for the most precision on the determination of $\delta$ for most cases. The higher energy is necessary for the longer baseline to get the appearance maximum to occur within the muon decay spectrum at that baseline. Depending on the specific details of the experimental configuration as well as progress in measuring the oscillation parameters at NOvA and T2K, SK and IceCube, as well as DUNE and HK, may lead to slight modifications on these optimal energies.
In fig. 2 we show the expected $1\sigma$ precision on $\delta$ as a function of CID and exposure for the FNAL-SURF setup for $\delta_{\text{true}}=0,-90^{\circ}$. We include the anticipated information from 10 years each of DUNE and HK. For DUNE we assume 480 kT-MW-year which corresponds to 5 years of each neutrino running and anti-neutrino with 1.2 MW proton beam and with a total fiducial volume of 40 kT of LAr. For HK we assume 190 kT water detector, 1.3 MW beam running for 10 years with $\nu:\bar{\nu}=1:3$ from Scott (2021). We simulate both experiments within the GLoBES framework.
@@
Since the beam at a neutrino factory consists of equal amounts of neutrinos and anti-neutrinos, there is an inherent large contribution to the final states coming from the other charge, i.e. the processes $\nu_{e}\to\nu_{\mu}$ and $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ lead to the same final state although with a different energy spectrum due to different initial spectra and different oscillation probabilities. Differentiating the events from the disappearance channel in the analysis of the disappearance channel has been identified as an important step towards a successful neutrino factory early on De Rujula et al. (1999). Here we study the impact of CID efficiency $\epsilon_{CID}$ in terms of either electron or muon charge identification. We define the CID parameter $\epsilon_{CID}$ as
@@
with the detection efficiency $\epsilon_{f}$ and $N_{\nu}$ is the number of (anti-)neutrinos of flavor $f$ we get after taking oscillations of the initial flux into account. DUNE could potentially differentiate a muon final state from an anti-muon final state with $\epsilon_{CID}\approx 72\%$ Ternes et al. (2019) from muon capture on argon. In addition, inelasticity can provide some CID information as well.
@@
Larger values of $\epsilon_{CID}$ could be achieved with magnetized detectors, an idea pursued by the INO collaboration. They estimate a CID efficiency of $\epsilon_{CID}\gtrsim 95\%$ for muons with GeV energy Ahmed et al. (2017). Electron CID at GeV energies has been studied in the scenario of a totally active scintillator detector in a magnetic field Bross et al. (2008); Asfandiyarov et al. (2014), and a magnetized LAr detector Rubbia (1977, 2001, 2004, 2009). Given that the final DUNE detector modules remain undesigned, a discussion of the possible benefit of working to improve CID is quite timely. In Huber and Schwetz (2008) it has been pointed out that CID can also be achieved statistically at some level, provided that the energy resolution is sufficiently good.

### v2

# A Modern Look at the Oscillation Physics Case for a Neutrino Factory

@@ Abstract
The next generation of neutrino oscillation experiments, JUNO, DUNE, and HK, are under construction now and will collect data over the next decade and beyond. As there are no approved plans to follow up this program with more advanced neutrino oscillation experiments, we consider here one option that had gained considerable interest more than a decade ago: a neutrino factory. Such an experiment uses stored muons in a racetrack configuration with extremely well characterized decays reducing systematic uncertainties and providing for more oscillation channels. Such a machine could also be one step towards a high energy muon collider program. We consider a long-baseline configuration to SURF using the DUNE far detectors or modifications thereof, and compare the expected sensitivities of the three-flavor oscillation parameters to the anticipated results from DUNE and HK. We show optimal beam configurations, the impact of charge identification, the role of statistics and systematics, and the expected precision to the relevant standard oscillation parameters in different DUNE vs. neutrino factory configurations.
1 Introduction
@@
The next generation of neutrino oscillation experiments, including JUNO JUNO:2022mxj , DUNE DUNE:2020ypp , and HK Hyper-KamiokandeProto-:2015xww , are expected to make significant headway in addressing three of the remaining known unknowns in particle physics. Notably, they will have excellent $\gg 5\sigma$ sensitivity to the mass ordering (MO) (the sign of $\Delta m^{2}_{31}$), they will have good $\theta_{23}$ octant determination capabilities provided that it isn’t too close to maximal, and both DUNE and HK will each individually provide $>5\sigma$ sensitivity to CP violation for significant regions of parameter space. In addition, they will provide some resolution on measuring $\delta$ (we use the standard parameterization of the PMNS Pontecorvo:1957cp ; Maki:1962mu matrix Denton:2020igp ). Atmospheric neutrino experiments such as HK, the IceCube-upgrade Ishihara:2019aao , and KM3NeT-ORCA KM3Net:2016zxf will also provide complementary information. For a recent overview of this picture, see Denton:2022een .
@@
While the outlook is hopefully rosy, it is important to carefully examine what, if anything, should come after these experiments in the neutrino oscillation program. Such an experiment would provide high precision measurements of all the oscillation parameters [^3] The solar parameters will not be well measured in long-baseline experiments Denton:2023zwa , but JUNO will provide high precision measurements of them JUNO:2022mxj and long-baseline experiments will probe the rest., test new physics scenarios in oscillations, and address any potential anomalies that arise in upcomming DUNE and HK measurements. Several experiments have been proposed with varying levels of detail. One is using a water-based liquid scintillator detector THEIA which combines the advantages of water and liquid scintillator, likely in place of one of the four LArTPC modules for DUNE Theia:2019non ; Theia:2022uyh . Another is an additional HK-like tank in Korea experiencing a different portion of the same beam that HK experiences to target the second oscillation maximum in appearance Hyper-Kamiokande:2016srs . There is also discussion about building a very large (much larger than HK) water Cherenkov detector in Sweden combined with additional beam upgrades of the ESS to create the ESSnuSB also to target the second oscillation maximum in ESSnuSB:2021azq ; Alekou:2022mav ; Abele:2022iml .
@@
In this article we discuss a different option, a neutrino factory. Neutrino factories were seriously considered as an option to probe CP violation in the first decade of this millennium largely due to the expectation of small $\theta_{13}<1^{\circ}$ which severely limits the possible size of the CP signal DeRujula:1998umv . In general, a neutrino factory is a muon storage ring in a racetrack configuration with long straight sections pointing towards the far detector and provides powerful beams of both muon and electron neutrinos of opposite charges Geer:1997iz ; Blondel:2000gj ; Blondel:2006su ; FernandezMartinez:2010zza . The initial spectrum of this decay is extremely well characterized unlike the initial spectrum from traditional fixed target neutrino oscillation experiments which primarily produce neutrinos from light meson decays with a range of energies. See Bogacz:2022xsj for a recent review of the history of neutrino factories as well as a discussion of the non-oscillation physics program at such a machine.
@@
A neutrino factory may also make sense as a first step on the path to a multi-TeV high energy muon collider to probe the electroweak sector, see e.g. Accettura:2023ked . This first step could be chronologically as a part of a technology demonstrator, a part of the accelerator chain, or both. In addition, the recent 2023 P5 report suggested consideration of neutrino physics to be done with a muon accelerator p5_2023 .
@@
In this paper we will discuss the oscillation physics at a neutrino factory considering the current knowledge of the oscillation parameters and address how a modern look at the oscillation physics case for a neutrino factory compares to previous studies. We then derive numerical sensitivities for the precision on $\delta$, $\theta_{23},\theta_{13}$ and $\Delta m_{31}^{2}$ for a potential setup of a neutrino factory for two different baselines and energies, before we conclude.
2 Neutrino Factory Setup
@@
  • The neutrino energies likely reach higher energies at a neutrino factory than a fixed target and the spectrum is different with part of it rising to the maximum, see fig. 1.
  • The composition and the expected energy of the neutrino beam is extremely well characterized and radiative corrections to muon decay are small Greub:1993kg ; Laing:2008zzb ; Tomalak:2021lif . Furthermore, no $\nu_{\tau}$ get produced in the source, allowing for cleaner searches for $\nu_{\tau}$ appearance channels.
  • The neutrino beam at a neutrino factory is composed of both $\nu_{\mu}$ and $\nu_{e}$ and of different charges allowing studies for channels not accessible in other controlled environments including $\nu_{e}\to\nu_{\mu}$ and $\nu_{e}\to\nu_{e}$; see Denton:2018cpu .
The first studies of a neutrino factory studies were focused on measuring the complex phase $\delta$ under the assumption of a very small $\theta_{13}$. These studies identified numerous interesting approximate degeneracies in the oscillation analysis Barger:2001yr ; Freund:2001ui ; Huber:2006wb . For small values of $\theta_{13}$, charge ID (CID) of muons or electrons was crucial to suppress the beam background for the appearance channels and to measure $\delta$. For this reason, many studies assumed a 100 kT magnetized iron detector IDS-NF:2011swj .
With the current and envisioned knowledge of all oscillation parameters and the present experimental landscape, the task of a future neutrino factory is no longer to discover the oscillation parameters, but rather to make precise measurements of the oscillation parameter and potentially resolve any discrepancies identified in previous measurements. Therefore, with a different physics goal and a more modern experimental landscape, the experimental requirements on a neutrino factory also necessarily evolve.
@@
We consider two possible neutrino factory scenarios: the neutrino source is Fermilab and, to make use of existing facilities, the far detector to be located at SURF, resulting in a baseline of 1284.9 km. Alternatively, we also consider the neutrino source at Brookhaven National Laboratory (BNL), to make use of the AGS/RHIC/EIC accelerator complex, which leads to baseline of 2542.3 km to SURF. [^4] See Kitano:2024kdv for the study of a neutrino factory based at J-PARC. We assume as matter density for both configurations 2.848 g/$\text{cm}^{3}$; the average matter density for the BNL configuration would actually be somewhat higher which would slightly enhance the differences between the two seen below. We assume that the far detector is a LArTPC configuration with a total fiducial target mass of 40 kT consistent with the envisioned DUNE detector. Based on previous exploratory studies of neutrino factories we assume $10^{21}$ muon (or anti-muon) decays per year IDS-NF:2011swj . This would require a comparable (to within $\mathcal{O}(1)$ factors) beam power as the 1.2 MW beam expected for DUNE.
@@
To simulate a neutrino factory we make use of DUNE’s far detector design DUNE:2021cuw in the GLoBES Huber:2004ka framework. We also consider oscillated tau neutrinos events with a tau neutrino CC cross section from psf2023008064 , although we treat the $\nu_{\tau}$ events as background only Donini:2010xk and don’t consider tau signal events for this three-flavor oscillation study due to the large uncertainties on the signal efficiency at LAr detectors. We have verified that including this channel would only marginally improve the sensitivity to measure the oscillation parameters because the $\nu_{\tau}$ cross section is significantly suppressed at these energies.
While we do not perform a full end-to-end simulation of DUNE or a neutrino factory as that is beyond the scope of this work and would require many detector and accelerator details that are still under development, we do implement uncertainties designed to match what is generally available. We use the files provided by DUNE in DUNE:2021cuw , which largely come from DUNE:2015lol and slightly modify them. For DUNE, we take the conservative estimates for the muon and electron neutrino uncertainties of 5% and 2%, respectively. The muon neutrino uncertainty taken to be larger than the electron neutrino uncertainty because it includes determinations of the beam flux, many cross section parameters, and the fiducial volume, consistent with the approach provided by DUNE. The electron neutrino uncertainty does include a contribution due to the translation of those measurements into the appropriate values for electron neutrinos. For the beam power we take 1.2 MW with 56% uptime as recommended [^5] Note that for the neutrino factory we consider only the total number of stored muons per year, which should thus have any uptime effects already accounted for..
For a neutrino factory going to the same far detector, we take the same approach with some improvements. We remove the uncertainty due to the detector volume as we only consider neutrino factory statistical tests simultaneously with DUNE measurements, so the same uncertainty should not be double counted. We also remove the flux uncertainty as it will be significantly reduced due to the well characterized beam from muon decay. Finally, we split the cross section model uncertainty equally across both flavors to find 2.5% uncertainty for both flavors because the statistics in the FD is comparable between both flavors.
For the backgrounds we keep with the recommendations of DUNE:2021cuw and 10% on the NC background, 20% on the tau background, and 2% on the Earth’s density.
3 Sensitivities
@@
For our simulations we consider priors on the oscillation parameters other than $\delta$ from Gonzalez-Garcia:2021dve and we fix the MO to the normal ordering (NO), as the MO as well as all oscillation parameters will be determined via multiple independent methods with future experiments. For comparison we also show the results for true inverted ordering IO in appendix B. We focus our sensitivity studies on $\delta$ as this is the least measured parameter in neutrino oscillations, but we also investigate a neutrino factory’s sensitivity to the other relevant oscillation parameters: $\Delta m^{2}_{31}$, $\theta_{23}$, and $\theta_{13}$, while long baseline experiments only have limited sensitivity to the solar parameters Denton:2023zwa . The precision on $\delta$ does provide our primary metric quantifying the ability of a neutrino factory to differentiate among differing measurements from upcoming experiments.
@@
Since the beam at a neutrino factory consists of equal amounts of neutrinos and anti-neutrinos, there is an inherent large contribution to the final states coming from the other charge, i.e. the processes $\nu_{e}\to\nu_{\mu}$ and $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ lead to the same final state although with a different energy spectrum due to different initial spectra and different oscillation probabilities, see appendix A. Differentiating the events from the disappearance channel in the analysis of the appearance channel has been identified as an important step towards a successful neutrino factory early on DeRujula:1998umv . Therefore we study the impact of CID efficiency $\epsilon_{CID}$ in terms of either electron or muon charge identification. We define the CID parameter $\epsilon_{CID}$ as
@@
with the detection efficiency $\epsilon_{f}$ and $N_{\nu}$ is the number of (anti-)neutrinos of flavor $f$ we get after taking oscillations of the initial flux into account. DUNE could potentially differentiate a muon final state from an anti-muon final state with $\epsilon_{CID}\approx 72\%$ Ternes:2019sak from muon capture on argon. In addition, inelasticity can provide some CID information as well.
@@
Larger values of $\epsilon_{CID}$ could be achieved with magnetized detectors, an idea pursued by the INO collaboration. They estimate a CID efficiency of $\epsilon_{CID}\gtrsim 95\%$ for muons with GeV energy ICAL:2015stm . Electron CID at GeV energies has been studied in the scenario of a totally active scintillator detector in a magnetic field Bross:2007ts ; Asfandiyarov:2014haa , and a magnetized LAr detector Rubbia:1977zz ; Rubbia:2001pk ; Rubbia:2004tz ; Rubbia:2009md . Given that the final DUNE detector modules remain undesigned, a discussion of the possible benefit of working to improve CID is quite timely. In Huber:2008yx it has been pointed out that CID can also be achieved statistically at some level, provided that the energy resolution is sufficiently good.
To compare our results for a NF and to get reliable conclusions about the potential state of oscillation physics with an inclusion of a NF after the next generation of LBL experiments we combine and compare the NF results to DUNE’s and HK’s forecasted results. For DUNE we assume 480 kT-MW-year which corresponds to 5 years of each neutrino running and anti-neutrino with 1.2 MW proton beam and with a total fiducial volume of 40 kT of LAr. For HK we assume 190 kT water detector, 1.3 MW beam running for 10 years with $\nu:\bar{\nu}=1:3$ from Scott:2020gng . We simulate both experiments within the GLoBES framework.
Figure 2: The expected $1\sigma$ precision on $\delta$ as a function of the muon energy for a total of 10 years at a NF (without inputs from DUNE/HK) assuming $10^{21}$ muon decays per year, for two different baselines and two different true values of $\delta$. The colors show different assumption of CID capabilities. For comparison, the horizontal lines indicate the expected sensitivity of 10 years of DUNE.
Figure 3: The expected $1\sigma$ precision on $\delta$ as a function of the running time of a NF for two different baselines and two different true values of $\delta$ and NF-FNAL energy of 5 GeV, NF-BNL energy of 8 GeV. We assume $10^{21}$ muon decays per year. We assume that the NF is combined with 10 years of DUNE. For comparison we show the expected sensitivity for 20 years of DUNE with the nominal systematic uncertainties, or assuming that electron and muon neutrino normalization uncertainties will be reduced to 1% in the final 10 years.
3.1 Sensitivity to $\delta$
In fig. 2 we show the sensitivity to $\delta^{\text{true}}=(-90^{\circ},~{}0)$ as a function of muon energy for a total of $10^{22}$ muon decays and a 40 kT FD. For the FNAL-SURF baseline we find that a muon energy of $E_{\mu}\simeq 5$ GeV with equal muon and anti-muon running time allows for the most precision on the determination of $\delta$ for most cases. For the BNL-SURF baseline we find a higher energy $E_{\mu}\simeq 8$ GeV is optimal, although higher energies have similar results. The higher energy is necessary for the longer baseline to get the appearance maximum to occur within the muon decay spectrum at that baseline. In general we find that the optimal region covers a range of energies as the neutrino flux at a NF has a broad spectrum. Assuming perfect CID abilities slightly increases the optimal energy range. For these optimal energies a NF can improve over 480 kT-MW-year of DUNE assuming $10^{22}$ total muon decays. At low muon energies a NF could be sensitive to the second oscillation maximum as demonstrated by the dips in the precision curve in fig. 2, however such low energy muons have a short lifetime which could lead to technical problems in realizing this setup.
Depending on the specific details of the experimental configuration as well as progress in measuring the $\delta$ and the other oscillation parameters may lead to slight modifications on these optimal energies. In the following we will fix the optimal muon energies to 5 GeV for the FNAL setup and 8 GeV for the BNL setup. [^6] In app. B we show the results for higher energies for the BNL setup as well.
In fig. 3 we show the $1\sigma$ precision on $\delta$ for $\delta^{\text{true}}=(-90^{\circ},~{}0)$ for the two setups for the optimal muon energies of 5 Gev and 8 GeV as a function of the total number of muon decays. We assume that the NF is combined with 480 kT-MW-year of DUNE. To improve the precision on $\delta$ independent of the true value of $\delta$ compared to 960 kT-MW-year of DUNE a total of $\gtrsim 11\times 10^{21}~{}\mu$ decays is required for the FNAL setup without CID but only $\gtrsim 4\times 10^{21}~{}\mu$ if CID can be realized. For the BNL setup even less muon decays are required, $\gtrsim 6\times 10^{21}$ decays without CID, $\gtrsim 3\times 10^{21}$ with CID. Even if the normalization uncertainty on the DUNE flux can be decreased to 1%, a NF can still lead to improvements on the precision over 480 kT-MW-year of DUNE with nominal uncertainties combined with 480 kT-MW-year with reduced uncertainties. In the following we fix the total number of muons to $10^{22}$ for a 40 kT fiducial mass FD. This could correspond to running of 10 years, equally divided between muons and anti-muons, with $10^{21}$ (anti-)muon per year.
To show the effect of CID in more detail we show in fig. 4 the expected $1\sigma$ precision on $\delta$ as a function of CID and exposure for the FNAL-SURF setup for $\delta_{\text{true}}=0,-90^{\circ}$. We include the anticipated information from 10 years each of DUNE and HK. Our results show that CID can improve the precision on $\delta$ by 15-20%. We have also studied the effect of assuming muon and electron CID at the same time and the results are similar. We find that electron CID has a bigger impact on the improvements in precision on $\delta$ than muon CID. This is likely due to the fact that the detected $\nu_{\mu}\to\nu_{e}$ spectrum is more similar to the detected $\bar{\nu}_{e}\to\bar{\nu}_{e}$ spectrum than the detected $\nu_{e}\to\nu_{\mu}$ spectrum to the detected $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ spectrum, see appendix A.
For the BNL baseline the improvement on the precision on $\delta$ does not significantly changes between 6-14 GeV, in particular if we assume CID. Constructing a neutrino factory with a higher muon energy could be preferential to avoid too many fast muon decays and for R&D studies of muon cooling technologies. Therefore we also study the precision on the oscillation parameters at these higher energies. The results can be found in the appendix B in tab. 3.